# Exploiting Unlabeled Data with Multiple Expert Teachers for Open Vocabulary Aerial Object Detection and Its Orientation Adaptation

### v1

@@ Abstract
In recent years, aerial object detection has been increasingly pivotal in various earth observation applications. However, current algorithms are limited to detecting a set of pre-defined object categories, demanding sufficient annotated training samples, and fail to detect novel object categories. In this paper, we put forth a novel formulation of the aerial object detection problem, namely open-vocabulary aerial object detection (OVAD), which can detect objects beyond training categories without costly collecting new labeled data. We propose CastDet, a CLIP-activated student-teacher detection framework that serves as the first OVAD detector specifically designed for the challenging aerial scenario, where objects often exhibit weak appearance features and arbitrary orientations. Our framework integrates a robust localization teacher along with several box selection strategies to generate high-quality proposals for novel objects. Additionally, the RemoteCLIP model is adopted as an omniscient teacher, which provides rich knowledge to enhance classification capabilities for novel categories. A dynamic label queue is devised to maintain high-quality pseudo-labels during training. By doing so, the proposed CastDet boosts not only novel object proposals but also classification. Furthermore, we extend our approach from horizontal OVAD to oriented OVAD with tailored algorithm designs to effectively manage bounding box representation and pseudo-label generation. Extensive experiments for both tasks on multiple existing aerial object detection datasets demonstrate the effectiveness of our approach. The code is available at https://github.com/lizzy8587/CastDet.
Index Terms:
Open-vocabulary aerial detection, Oriented object detection, Semi-supervised learning, Multi-modal.
I Introduction
@@
Object detection in aerial images aims at localizing and categorizing objects on the earth’s surface, which is fundamental to a wide range of remote sensing applications, such as urban planning, environmental surveillance, and disaster response [1, 2, 3]. Despite significant advancements in aerial object detectors powered by deep learning [4, 5, 6, 7, 8, 9], they are still weak at detecting objects beyond the training categories.
@@
To expand the object detectors to novel categories, a naive idea is to collect and annotate large-scale aerial images with rich object categories. However, obtaining accurate yet sufficient annotations is time- and labor-expensive, even requiring human experts involvement. As a result, despite extensive efforts paid in the collection of existing aerial object detection datasets [10, 11, 12, 13, 14, 15], they are still much smaller and less diverse than natural image datasets [16, 17, 18], as depicted in Fig. 1. It hinders the detector scaling up in open-world scenarios. Therefore, this paper advocates more flexible object detectors capable of detecting novel object categories unseen during training, referred to as open vocabulary object detection (OVD). This allows us to characterize new objects that emerged in the earth observation data without extra annotations in open scenarios.
Figure 1: Comparisons of the amount of categories and images of 19 common aerial and natural datasets. Challenge 1: Aerial datasets are much smaller in size and category vocabularies compared with natural image datasets.
@@
Drawing inspiration from the recent success of OVD in natural images [19, 20, 21], we intend to explore ways in the challenging task, namely open vocabulary aerial object detection (OVAD), where objects often vary widely in scale, orientation, and exhibit weak feature appearance [22]. Unlike natural images, which often feature clear contours and textures, allowing class-agnostic region proposal networks (RPNs) to effectively generalize to novel categories [23, 24]; aerial images captured from overhead perspectives usually display weak surface features, where objects may blend into the background. This complicates object detection, for instance, AIRPORT may be confused with HIGHWAY, and thus be wrongly classified as background, as illustrated in Fig. 2(b). Consequently, the recall of novel categories in aerial images is significantly lower than that in natural images, as shown in Fig. 2(c). In addition, objects in aerial images exhibit great orientation diversity. Using horizontal bounding boxes for these rotated objects could bring in substantial background clutter, leading to inaccurate localization. However, current open-vocabulary detectors primarily support horizontal detection, which limits their effectiveness in aerial images. These factors, on the one hand, pose challenges to directly apply current OVD methods for natural images to aerial images; and, on the other hand, spur us to develop extensible object detectors for aerial images, covering more object categories without extra annotation.
@@
Figure 2: Challenge 2: The recall of novel objects in aerial images is much lower than that in natural images due to the highly complex backgrounds. (a)(b) Aerial images from DIOR [11]. Objects in aerial images exhibit background interference. (c) Class-agnostic RPN recall statistics of novel categories in the natural dataset COCO [16] and the aerial dataset VisDroneZSD [25] (i.e., 77% v.s. 48%).
@@
To address these challenges, we propose a simple yet effective open-vocabulary aerial detection framework, namely CastDet, a CLIP-activated student-teacher detector. It comprises three key models: a student model and two teacher models for localization and classification, respectively. The student model is responsible for training both horizontal and oriented open-vocabulary detectors. The former adopts methods such as Faster R-CNN [26], and the later exploits the Oriented R-CNN [27]. The localization teacher model focuses on discovering and localizing potential objects, while the other teacher model provides classification guidance for novel categories. We jointly train the student model and teacher models in a self-learning manner, where the student is optimized with pseudo-labels generated by the teachers, and the localization teacher is updated from the continually learned student model. As exemplified by the literature [28], the slowly progressing localization teacher model acts as an ensemble of the previous student models, exhibiting superior capabilities that enable it to generate more reliable proposals. These proposals are then filtered based on their confidence scores, such as scale jittering variance, before being presented to the external teacher.
@@
While the student-teacher learning paradigm is powerful for enhancing the capabilities of the student model under the guidance of the teacher model [29, 30, 31, 28, 32], they operate in the closed-set setting. This leads to poor performance in discovering and recognizing novel object categories that are not encountered in the training data. To overcome this limitation, RemoterCLIP [33] that has been pre-trained on massive remote sensing image-text pairs and equips with strong zero-shot classification capabilities is exploited as an external teacher model. Thus, the learning process could benefit from the extra knowledge dedicated to interpreting aerial images brought by the RemoteCLIP. In addition to involving the RemoteCLIP, maintaining high-quality pseudo-labels is necessary. To this end, after predicting the category labels for the pseudo proposals, a dynamic label queue is proposed to store and iteratively update the pseudo-labels.
@@
Unlike previous CLIP-based methods [34, 35, 23] that directly transfer knowledge from CLIP for open-vocabulary recognition, our CLIP-activated student-teacher self-learning framework incorporates high-confidence knowledge from RemoteCLIP as an incentive to guide both the student and localization teacher. This self-learning mechanism facilitates a “flywheel effect” wherein the external teacher transfers knowledge to strengthen the localization teacher to identify potential regions of novel objects, thereby enhancing pseudo box quality. In turn, the refined pseudo boxes enable the external teacher to produce more precise pseudo-labels. Through our student-teacher self-learning process, the detection model can be progressively updated to localize and recognize continuously expanded object category vocabularies.
@@
The preliminary content of this paper has partially appeared in the conference version, with the detector named CastDet [36] (ECCV 2024). The overall contributions of this extended journal version are as follows [^1] This journal version extends the previous conference version (ECCV 2024) [36], particularly in the following aspects: 1) We rewrite the full paper with comprehensive discussions and more details; 2) The proposed CastDet framework is now extended to support open-vocabulary oriented detection, addressing the specific challenges of detecting objects with arbitrary orientations in aerial scenes; 3) We introduce two new oriented box selection strategies: box scale jittering variance and angle jittering variance, which take both the scale and orientation of objects into account to filter robust pseudo boxes; 4) We integrate multiple open-vocabulary oriented detection algorithms into a unified code repository, including Oriented CastDet, Oriented ViLD, Oriented GLIP and Oriented GroundingDINO; 5) We conduct more ablation experiments to demonstrate the effectiveness of the proposed techniques; 6) We build a more comprehensive benchmark for the open-vocabulary aerial detection task, including both horizontal object detection and oriented object detection.:
@@
    We introduce the CastDet, a flexible open-vocabulary detection framework incorporating the student-(multi)teacher self-learning paradigm. The proposed framework enables the detection model to progressively expand its object vocabulary and achieve accurate localization of novel objects using horizontal or oriented boxes without extra annotation efforts.
@@
    We devise a dynamic label queue for storing and incrementally updating high-quality pseudo-labels generated by the teachers. By this way, richer and more accurate labels could be dynamically maintained.
@@
II Related work
II-A Fully-supervised Object Detection
@@
Object detection aims to predict the bounding box coordinates and corresponding category labels for objects of interest within an image. Generally, existing object detection methods can be divided into two-stage and one-stage methods. Two-stage methods [37, 38, 26, 39, 40] such as Faster R-CNN [26] initially employ a region proposal network to generate potential target regions, then classify and refine these proposals. One-stage methods model the object localization and categorization as an end-to-end problem, which is to be resolved through a single network. Classical methods include anchor-based ones, such as YOLO [41], SSD [42], RetinaNet, [43], and anchor-free methods, e.g., CornerNet [44], FCOS [45], CenterNet [46].
@@
Inspired by the advanced progress of object detection methods [26, 43, 41, 47, 22] dedicated to natural images, efforts have been paid to adapt these methods to aerial images. Specifically, several challenges remains to be tackled in aerial images, including object orientation [5, 27, 48, 6, 7], scale variation [49, 50], and dense objects [51, 8, 9, 52]. Early works [4, 53, 54, 55] mainly focus on horizontal detection using classical deep learning methods to detect aerial objects with horizontal bounding boxes. However, horizontal bounding box could cause inaccurate location of aerial objects with arbitrary orientations. Therefore, methods that could represent objects by oriented bounding boxes [51, 56] have attracted widespread attention. One focus of research is on optimizing the representation of oriented objects, which can be achieved by converting the angle regression task into an angle classification task [57, 58] or by modeling oriented objects using a Gaussian distribution [59, 60, 61]. In parallel, other studies emphasize learning robust alignment feature for oriented objects, exemplified by approaches such as SCRDet [8], R3Det [6], S2A-Net [62], and ReDet [48].
II-B Semi-supervised Object Detection
Semi-supervised learning (SSL) incorporates external unlabeled data to improve performance through various techniques, including data augmentation [63, 64, 65], pseudo-labeling [66, 67, 68], and consistency training [69, 70, 71, 72]. CSD [29] trains the detector to match the predictions between the original and horizontal flipped images. STAC [30] generates high confidence offline pseudo-labels from unlabeled images, and updates the detector through consistency training with strong augmentations. Instant-Teaching [31] updates pseudo-labels during the model training thus improving the quality of pseudo annotations. Furthermore, several works incorporate an Exponential Moving Average (EMA) strategy inspired by the Mean Teacher [70] to continuously update the teacher model [28, 32, 73, 74]. Among these, Unbiased Teacher [28] further addresses pseudo-label bias through focal loss; Soft-Teacher [32] employs soft labels weighted by confidence scores, and a box jittering strategy to select reliable pseudo boxes for training the box regression branch.
Recent studies have explored semi-supervised oriented object detection [75, 76, 77, 78, 79], which presents additional challenges due to the arbitrary orientations of objects. SOOD [75, 80] proposes rotation-aware adaptive weighting loss and global consistency loss for semi-supervised oriented object detection. S2O-Det [79] introduces coarse-to-fine sample mining strategy for selecting discriminative pseudo-labels. Focal Teacher [77] employs a global focal learning method to generate high-quality pseudo-labels without relying on artificial priors. Despite these advances, existing methods still face certain limitations. Current aerial object detectors operate under a closed-set assumption, where all categories of interest are annotated before training. Detectors trained in such a closed-set manner struggle to identify novel categories, limiting their adaptability to open world.
II-C Open-vocabulary Object Detection
@@
Open-vocabulary object detection (OVD) aims to detect objects beyond a close-set vocabularies. OVR-CNN [81] pioneered this approach by combining bounding box annotations for a subset of categories with a corpus of image-caption pairs to extend the vocabulary of detectable objects. Leveraging the zero-shot capabilities of pre-trained Vision-Language Models (VLMs), such as CLIP [82], recent OVD methods transfer knowledge from pre-trained VLMs through prompt learning [83, 84, 85] or region-level fine-tuning [34, 35, 24, 86] to achieve flexible and versatile detection of extensible object categories. For instance, ViLD [34] transfers knowledge from a pre-trained VLM to enrich a two-stage detector through both vision and language knowledge distillation. RegionCLIP[35] aligns visual representations at the region level with text descriptions. Detic [24] enhances the vocabularies by incorporating image classification data, expanding the scope of detectable objects to tens of thousands. PromptDet [84] and DetPro [83] leverage prompt learning techniques to enhance the alignment between region visual feature and semantic feature. Additionally, some works unify object detection and phrase grounding tasks to pre-train a detector capable of open-vocabulary detection [19, 87, 21]. For example, GLIP [19] integrates image-language pre-training to align visual and textual data, improving object detection across diverse categories using text prompts. GroundingDINO [21] combines the DINO detector with grounded pre-training, enhancing open-set object detection by tight modality fusion between vision and language with large-scale datasets. The success of these approaches relies on two critical factors: 1) effective object proposal generation beyond training categories; and 2) large-scale image-text datasets that enable robust open-vocabulary classification capabilities.
@@
The limited size of existing aerial image datasets, along with their distinct visual characteristics compared to natural images, leads to suboptimal generalization in region proposal generation for diverse categories. Consequently, open-vocabulary detection methods developed for natural images perform poorly on aerial images. To address the gap in open-vocabulary aerial detection, we introduce CastDet [36], the first approach designed specifically for this task, which tackles the unique challenges posed by aerial imagery through a student-multi-teacher self-learning paradigm. Following CastDet, a few studies further explore OVAD. For example, LAE-DINO [88] proposes a DINO-based detector with dynamic vocabulary construction and visual-guided text prompt learning to improve detection across diverse aerial categories. OVA-DETR [89] enhances feature extraction by leveraging image-text alignment with region-text contrastive loss and Bidirectional Vision-Language Fusion, particularly for small objects. However, none of these methods address the challenge of oriented object detection in aerial scenes. This paper aims to fill this blank and set a foundation for future research.
III Method

### v2

@@ Abstract
In recent years, aerial object detection has been increasingly pivotal in various earth observation applications. However, current algorithms are limited to detecting pre-annotated categories. In this paper, we put forth a novel formulation of the aerial object detection problem, namely open-vocabulary aerial object detection (OVAD), which can detect objects beyond training categories without costly collecting new labeled data. We propose CastDet, a CLIP-activated student-teacher detection framework that serves as the first OVAD detector specifically designed for the challenging aerial scenario, where objects exhibit weak appearance features and arbitrary orientations. Our framework integrates a robust localization teacher along with several box selection strategies to generate high-quality proposals for novel objects. Additionally, the RemoteCLIP model is adopted as an omniscient teacher, which provides rich knowledge to enhance classification capabilities for novel categories. A dynamic label queue is devised to maintain high-quality pseudo-labels during training. By doing so, the proposed CastDet boosts not only novel object proposals but also classification. Furthermore, we extend our approach from horizontal to oriented OVAD with tailored algorithm designs for bounding box representation and pseudo-label generation. Extensive experiments for both tasks on multiple aerial datasets demonstrate the effectiveness of our approach.
[…]
1 Introduction
@@
Object detection in aerial images aims at localizing and categorizing objects on the earth’s surface, which is fundamental to a wide range of remote sensing applications, such as urban planning, environmental surveillance, and disaster response [zhao2003car, reilly2010detection, sadgrove2018real]. Despite significant advancements in aerial object detectors powered by deep learning [7926624, ding2019learning, yang2021r3det, qian2022rsdet++, yang2019scrdet, yang2019clustered], they are still weak at detecting objects beyond the training categories.
@@
To expand the object detectors to novel categories, a naive idea is to collect and annotate large-scale aerial images with rich object categories. However, obtaining accurate yet sufficient annotations is time- and labor-expensive, even requiring human experts involvement. As a result, despite extensive efforts paid in the collection of existing aerial object detection datasets [nwpu_cheng2016learning, dior_li2020object, DOTA8578516, levir_zou2017random, visdrone_zhu2021detection, li2024star], they are still much smaller and less diverse than natural image datasets [lin2014microsoft, gupta2019lvis, 5206848imagenet], as depicted in Fig. 2. It hinders the scalability of detectors in open-world scenarios. Therefore, this paper advocates more flexible object detectors capable of detecting novel object categories unseen during training, referred to as open vocabulary object detection (OVD). This allows us to characterize new objects that emerged in the earth observation data without extra annotations in open scenarios.
@@
Drawing inspiration from the recent success of OVD in natural images [9879567glip, yao2022detclip, liu2023groundingdino], we intend to explore ways in the challenging task, namely open vocabulary aerial object detection (OVAD), where objects often vary widely in scale, orientation, and exhibit weak feature appearance [Zhang2023]. Unlike natural images, which often feature clear contours and textures, allowing class-agnostic region proposal networks (RPNs) to effectively generalize to novel categories [chen2023ovarnet, zhou2022detic]; aerial images captured from overhead perspectives usually display weak surface features, where objects may blend into the background, as shown in Fig. 3(a). This complicates object detection, for instance, AIRPORT may be confused with HIGHWAY, and thus be wrongly classified as background, as illustrated in Fig. 3(b). Consequently, the recall of novel categories in aerial images is significantly lower than that in natural images, as demonstrated in Fig. 2. In addition, objects in aerial images exhibit great orientation diversity. Using horizontal bounding boxes for these rotated objects could bring in substantial background clutter, leading to inaccurate localization, as shown in Fig. 3(c). However, current open-vocabulary detectors primarily support horizontal detection, which limits their effectiveness in aerial images. These factors, on the one hand, pose challenges to directly apply current OVD methods for natural images to aerial images; and, on the other hand, spur us to develop extensible object detectors for aerial images, covering more object categories without extra annotation.
Figure 1: Comparisons of the amount of categories and images of 19 common aerial and natural image datasets. Aerial datasets are much smaller in size and category vocabularies compared with natural image datasets.
@@
Figure 2: Class-agnostic RPN recall statistics of novel categories in the natural dataset COCO [lin2014microsoft] and the aerial dataset VisDroneZSD [VisDrone2023] (i.e., 77% v.s. 48%). The recall of novel objects in aerial images is much lower than that in natural images due to the highly complex backgrounds.
Figure 3: Aerial images from DIOR [dior_li2020object]. (a)(b) Objects in aerial images exhibit background interference. The red box represents the ground truth, while the yellow box highlights the areas of interference. (c) Inaccurate localization with horizontal bounding box (HBox), causing significant background clutter in aerial images compared to oriented bounding box (RBox).
@@
To address these challenges, we propose a simple yet effective open-vocabulary aerial detection framework, namely CastDet, a CLIP-activated student-teacher detector. It comprises three key models: a student model and two teacher models for localization and classification, respectively. The student model is responsible for training both horizontal and oriented open-vocabulary detectors. The former adopts methods such as Faster R-CNN [ren2015faster], and the later exploits the Oriented R-CNN [orientedrcnn_Xie_2021_ICCV]. The localization teacher model focuses on discovering and localizing potential objects, while the other teacher model provides classification guidance for novel categories. We jointly train the student model and teacher models in a self-learning manner, where the student model is optimized with pseudo-labels generated by the teachers, and the localization teacher model is updated from the continually learned student model. As exemplified by the literature [liu2021unbiased], the slowly progressing localization teacher model acts as an ensemble of the previous student models, exhibiting superior capabilities that enable it to generate more reliable proposals. These proposals are then filtered based on their confidence scores, such as scale jittering variance, before being presented to the external teacher model.
@@
While the student-teacher learning paradigm is powerful for enhancing the capabilities of the student model under the guidance of the teacher model [jeong2019consistency, sohn2020simple, zhou2021instant, liu2021unbiased, softteacher_xu2021end], they operate in the closed-set setting. This leads to poor performance in discovering and recognizing novel object categories that are not encountered in the training data. To overcome this limitation, RemoteCLIP [liu2023remoteclip] that has been pre-trained on massive remote sensing image-text pairs and equips with strong zero-shot classification capabilities is exploited as an external teacher model. Thus, the learning process could benefit from the extra knowledge dedicated to interpreting aerial images brought by the RemoteCLIP. In addition to involving the RemoteCLIP, maintaining high-quality pseudo-labels is necessary. To this end, after predicting the category labels for the pseudo proposals, a dynamic label queue is proposed to store and iteratively update the pseudo-labels.
@@
Unlike previous CLIP-based methods [gu2021open, zhong2022regionclip, chen2023ovarnet] that directly transfer knowledge from CLIP for open-vocabulary recognition, our CLIP-activated student-teacher self-learning framework incorporates high-confidence knowledge from RemoteCLIP as an incentive to guide both the student model and localization teacher model. This self-learning mechanism facilitates a “flywheel effect” wherein the external teacher model transfers knowledge to strengthen the localization teacher model to identify potential regions of novel objects, thereby enhancing pseudo box quality. In turn, the refined pseudo boxes enable the external teacher model to produce more precise pseudo-labels. Through our student-teacher self-learning process, the detection model can be progressively updated to localize and recognize continuously expanded object category vocabularies.
@@
The preliminary content of this paper has partially appeared in the conference version, with the detector named CastDet [castdet] (ECCV 2024). The overall contributions of this extended journal version are as follows [^1] This journal version extends the previous conference version (ECCV 2024) [castdet], particularly in the following aspects: 1) We rewrite the full paper with comprehensive discussions and more details; 2) The proposed CastDet framework is now extended to support open-vocabulary oriented detection, addressing the specific challenges of detecting objects with arbitrary orientations in aerial scenes; 3) We introduce two new oriented box selection strategies: box scale jittering variance and angle jittering variance, which take both the scale and orientation of objects into account to filter robust pseudo boxes; 4) We integrate multiple open-vocabulary oriented detection algorithms into a unified code repository, including Oriented CastDet, Oriented ViLD, Oriented GLIP and Oriented GroundingDINO; 5) We conduct more ablation experiments to demonstrate the effectiveness of the proposed techniques; 6) We build a more comprehensive benchmark for the open-vocabulary aerial detection task, including both horizontal object detection and oriented object detection.:
@@
    We introduce the CastDet, a flexible open-vocabulary detection framework incorporating the student-(multi)teacher self-learning paradigm. The proposed framework enables the detection model to progressively expand its object vocabulary and achieves accurate localization of novel objects using horizontal or oriented boxes without extra annotation efforts.
@@
    We devise a dynamic label queue for storing and incrementally updating high-quality pseudo-labels generated by the teacher models. By this way, richer and more accurate labels could be dynamically maintained.
@@
2 Related work
2.1 Close-set Object Detection
@@
Object detection aims to predict the bounding box coordinates and corresponding category labels for objects of interest within an image. Generally, existing object detection methods can be divided into two-stage and one-stage methods. Two-stage methods [rcnn_girshick2014rich, fastrcnn_girshick2015fast, ren2015faster, cascadercnn_cai2018cascade, maskrcnn_he2017mask] such as Faster R-CNN [ren2015faster] initially employ a region proposal network to generate potential target regions, then classify and refine these proposals. One-stage methods model the object localization and categorization as an end-to-end problem, which is to be resolved through a single network. Classical methods include anchor-based ones, such as YOLO [yolo_redmon2016you], SSD [ssd_liu2016ssd], RetinaNet, [retina_lin2017focal], and anchor-free methods, e.g., CornerNet [law2018cornernet], FCOS [tian2019fcos], CenterNet [zhou2019objects].
@@
Inspired by the advanced progress of object detection methods [ren2015faster, retina_lin2017focal, yolo_redmon2016you, carion2020end, Zhang2023] dedicated to natural images, efforts have been paid to adapt these methods to aerial images. Specifically, several challenges remains to be tackled in aerial images, including object orientation [ding2019learning, orientedrcnn_Xie_2021_ICCV, han2021ReDet, yang2021r3det, qian2022rsdet++], scale variation [10509806_glh_bridge, ZHENG20201_HyNet], and dense objects [Yang2018AutomaticSD, yang2019scrdet, yang2019clustered, xu2022rfla]. Early works [7926624, zhang2016s, kang2017modified, liu2017sar] mainly focus on horizontal detection using classical deep learning methods to detect aerial objects with horizontal bounding boxes. However, horizontal bounding box could cause inaccurate localization of aerial objects with arbitrary orientations. Therefore, methods that could represent objects by oriented bounding boxes [Yang2018AutomaticSD, azimi2018towards, lee2023hausdorff] have attracted widespread attention. One focus of research is on optimizing the representation of oriented objects, which can be achieved by converting the angle regression task into an angle classification task [yang2022arbitrary, yang2021dense] or by modeling oriented objects using a Gaussian distribution [yang2021rethinking, yang2021learning, yang2022detecting]. In parallel, other studies emphasize learning robust alignment feature for oriented objects, exemplified by approaches such as SCRDet [yang2019scrdet], R3Det [yang2021r3det], S2A-Net [han2021align], and ReDet [han2021ReDet].
2.2 Open-vocabulary Object Detection
@@
Open-vocabulary object detection (OVD) aims to detect objects beyond the categories known in the training phase. OVR-CNN [zareian2021open] pioneered this approach by combining bounding box annotations for a subset of categories with a corpus of image-caption pairs to extend the vocabulary of detectable objects. Leveraging the outperforming zero-shot capabilities of pre-trained Vision-Language Models (VLMs), such as CLIP [CLIP_radford2021learning], recent OVD methods success in achieving flexible and versatile detection with extensible object categories by means of prompt learning [du2022learning, feng2022promptdet, zang2022open] or region-level fine-tuning [gu2021open, zhong2022regionclip, zhou2022detic, wu2023cora]. For instance, ViLD [gu2021open] transfers knowledge from the CLIP to enrich a two-stage detector via both vision and language knowledge distillation. RegionCLIP[zhong2022regionclip] aligns visual representations at the region level with text descriptions. Detic [zhou2022detic] enriches the object vocabulary by incorporating image classification data to expand the scope of detectable objects to tens of thousands. PromptDet [feng2022promptdet] and DetPro [du2022learning] take prompt learning techniques to enhance the alignment between visual feature and semantic feature at the region level. Additionally, some works [9879567glip, zhang2022glipv2, liu2023groundingdino] pre-train the detector capable of open-vocabulary detection with unified object detection and phrase grounding objectives. For example, GLIP [9879567glip] integrates image-language pre-training to align visual and textual data, improving object detection across diverse categories using text prompts. GroundingDINO [liu2023groundingdino] combines the DINO detector with grounded pre-training on large-scale datasets to enhance open-set object detection by the tight fusion between vision and language. The success of these approaches relies on two critical factors: 1) effective object proposal generation beyond training categories; and 2) large-scale image-text datasets that enable robust open-vocabulary classification capabilities.
@@
The limited size of existing aerial image datasets, along with their distinct visual characteristics compared to natural images, leads to unsatisfying generalization in region proposal generation for diverse categories. Consequently, open-vocabulary detection methods developed for natural images perform poorly on aerial images. To address the issue in open-vocabulary aerial detection, we introduce the CastDet [castdet], the first approach dedicated to tackle the unique challenges posed by aerial imagery through a student-multi-teacher self-learning paradigm. In addition, a few studies have been exploring the OVAD following the CastDet. For example, LAE-DINO [pan2024locate] proposes a DINO-based detector with dynamic vocabulary construction and visual-guided text prompt learning to improve detection across diverse aerial categories. OVA-DETR [wei2024ovadetropenvocabularyaerial] enhances feature extraction by leveraging image-text alignment with region-text contrastive loss and Bidirectional Vision-Language Fusion, particularly for small objects. However, none of these methods address the challenge of oriented object detection in aerial scenes. To this end, another goal of this paper is to fill this blank and set a foundation for future research.
3 Methodology